OpenStudy (anonymous):

Write 1/5 ,+ 1/6, + 1/7 in sigma notation

zepdrix (zepdrix):

\[\large \sum_{n=?}^? \frac{1}{n}\] Hmm it looks like we're just adding a bunch of \(\huge \frac{1}{n}\)'s. :) Can you determine where it will start and end?

OpenStudy (anonymous):

7 and 5?

OpenStudy (anonymous):

will it be that? @zepdrix

zepdrix (zepdrix):

Sorry the website is going reallllyyy slow :( Yessss it appears we're starting at n=5, giving us \(\huge \frac{1}{5}\) for our first term. Ending at n=7. Good good. c:

OpenStudy (anonymous):

thanks